# Wireless Network Topology Inference: A Markov Chains Approach

### v1

@@ 5 Experimental results
In Fig. 1 we present experimental results for a network topology corresponding to an undirected cycle of $6$ nodes. We can see that both transfer entropy and our methodology works well in recovering links in the network. In Fig. 1(d) we stud the effect of a varying transmission length on network estimation. In particular, we record the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window. More precisely, we say that a link $\{i,j\}$ has been correctly identified by our methodology (resp. transfer entropy) if it appears in the top $m$ entries of the upper triangular part of the $\hat{L}_{\text{sym}}$ (resp. transfer entropy) matrix, where $m$ in the number of links in the network. We observe that a shorter transmission window makes estimation harder since it implies a larger number of nodes will be concurrently transmitting (i.e., a larger parameter $k$ in 1). We notice how our procedure slightly improves upon transfer entropy when the network is congested.
Further experimental results are reported in Figs. 2, 3, 4 and 4. We highlight our methodology consistently outperforms transfer entropy whenever the network is congested. Transfer entropy and our estimate appear very similar when $T\geq 15s$.
@@
Figure 1: Experimental results for a cyclic network topology of six nodes. Figure (a) displays the adjacency matrix of the network. Figure (b) the estimated matrix $\hat{L}_{\text{sym}}$ for a transmission window length $T=15s$. Figure (c) the transfer entropy matrix ($T=15s$). Figure (d) represents the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window. All results are averaged over $50$ independent simulations.
@@
Figure 2: Experimental results for a network topology of eight nodes.
@@
Figure 3: Experimental results for a network topology where nodes are arranged in a three-by-three grid.
@@
Figure 4: Experimental results for a network topology where nodes are arranged in a four-by-two grid.
@@
Figure 5: Experimental results for a network topology of twenty nodes arranged in a path.
@@ 6 Conclusion and future work
In this work, we introduced a novel technique for estimating the topological structure of wireless networks. We demonstrated that our approach provably recovers information about network topology under a simplified Markovian model of network dynamics. Additionally, our experimental results showed that the proposed method is competitive with transfer entropy in realistic network simulations, consistently outperforming it under conditions of network congestion.
@@
Finally, measures such as transfer entropy and Granger causality are often used as features in more sophisticated estimation procedures. For instance, Testi and Giorgetti [TG20] utilise these measures to train a neural network model. It would be interesting to explore the integration of our techniques into machine learning frameworks for network inference.
@@ 7 Proofs
We will show that, for $T$ large enough, $M(u,v)\approx P(u,v)-(k-1)\pi(v)$ and $N(v)\approx T\cdot k\cdot\pi(v)$. Therefore, by computing $M$ and $N$, we are able to approximate $P$. We start by showing that $N(v)$ is approximately a multiple of the stationary distribution at $v$. For technical reason, we also need a bound on the following quantities ($u,v\in V$):
$\widetilde{N}(u,v)=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\},$ (2)
$Q(u,v)=\sum_{t=0}^{T-1}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(i^{\prime})}=v\}.$ (3)
We will use the following recent matrix Bernstein inequality for Markov chains [matrixMarkovBernstein].
@@
Let $\epsilon\in(0,1/2),\delta\in(0,1)$. Let $T>Ck^{2}(\pi^{\star}/\pi_{\star})\ln(kn/\delta)/(\epsilon^{2}\lambda(P)\pi_{\star})$ for some large enough universal constant $C>0$. Let $D_{N}$ be the $n\times n$ diagonal matrix such that $D_{N}(u,u)=N(u)$. Then, with probability at least $1-\delta$, for all $u,v\in V$, the following holds:
@@
$\left\|Q-Tk(k-1)\pi\pi^{T}-D_{N}\right\|\leq Tk^{2}\pi^{\star}$
@@
$\mathop{\mathbb{E}}[N(v)]$ $=\sum_{i=1}^{k}\sum_{t=0}^{T-1}\mathop{\mathbb{P}}[X^{(j)}_{t}=v]=kT\pi(v),$
@@
$\mathop{\mathbb{E}}[\widetilde{N}(u,v)]$ $=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathop{\mathbb{E}}[\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}]$
@@
Notice that $\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]=\sum_{i\neq j}\sum_{t=0}^{T-1}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})$. Clearly, $\|g_{ij}((u,v))\|\leq 2$ for all $(u,v)\in V\times V$. Furthermore, simple calculations show that
$\left\|\mathop{\mathbb{E}}_{#1}[g_{ij}((u,v))^{2}]\right\|\leq 2\pi^{\star}.$
Therefore, with probability at least $\delta/k^{2}$, Theorem 2 implies that
$\left\|\sum_{t=0}^{T-1}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})\right\|\leq Tk\pi_{\star}.$
A union bound over all $1\leq i\neq j\leq k$ yields the second statement.
Finally, the third inequality follows simply from the first two by noticing that $Q(u,v)=\widetilde{N}(u,v)+D_{N}(u,v)$, ∎
We now prove concentration for $M$. We apply similar techniques as in Kontorovich and Wolfer [Wolf19]. We will need the following concentration inequality for matrix martingales, which is due to Tropp [Tropp].
@@
where with $\mathop{\mathbb{E}}_{#1}[Y_{k}]$ we denote the conditional expectation of $Y_{k}$ wrt “the past”.
@@
$\mathop{\mathbb{E}}_{#1}[\mathbbm{1}\{X_{t-1}^{(i)}=u\}\mathbbm{1}\{X_{t}^{(j)}=v\}]=\mathbbm{1}\{X_{t-1}^{(i)}=u\}\sum_{w\in V}\mathbbm{1}\{X_{t-1}^{(j)}=w\}P(w,v)$
@@
Let $\widetilde{N}(u,v)=\sum_{t=0}^{T-1}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}$. Then, for any $u,v\in V$,
@@
$\|\widetilde{N}P-k(k-1)T\pi\pi^{T}\|$ $<\|\mathop{\mathbb{E}}[\widetilde{N}]P-k(k-1)T\pi\pi^{T}\|+\|(\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}])P\|$
@@
$\leq\max_{v\in V}\left\{|\sum_{i\colon X^{(i)}_{t}=v}(1-P(X^{(i)}_{t-1},v))|+k\cdot\sum_{i=1}^{k}|\sum_{j\colon X^{(j)}_{t}=v}(1-P(X^{(j)}_{t-1},v))|\right\}$
@@
$\sum_{v\in V}\left|\left(\mathbbm{1}\{X_{t}^{(i)}=v\}-P(u,v)\right)\right|$ $=1-P(u,X_{t}^{(i)})+\sum_{w\neq X_{t}^{(i)}}P(u,w)$
$=2(1-P(u,X_{t}^{(i)})).$
@@
$\leq\max_{u\in V}\left\{2\sum_{i=1}^{k}(1-P(u,X_{t}^{(i)}))+\sum_{i=1}^{k}\cdot|\sum_{j\neq i}2(1-P(X^{(j)}_{t-1},v))|\right\}$
@@
$\|Y_{t}\|\leq\sqrt{\|Y_{t}\|_{1},\|Y_{t}\|_{\infty}}\leq\sqrt{2}k^{2}.$
@@
$\qquad\qquad\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)$
@@
$\qquad\left(\mathbbm{1}\{X_{j}^{(i)}=v\}-P(w,v)\right)\Bigr{]}$
@@
Moreover, by Lemma 3, with high probability, we have that $N(u)\leq 2kT\pi(u)$ for any $u\in V$. Therefore,
@@
We have proved that $\|N-(D_{N}+\widetilde{N})P\|=\|D_{N}(M-P-\widetilde{M}P))\|\leq(1/\sqrt{2})\cdot Tk\pi_{\star}\epsilon$ with high probability. By (5), we have that $\|D_{N}(M-P-(k-1)\Pi)\|\leq(5/6)\cdot Tk\pi_{\star}\epsilon$ with high probability. Since $\min_{u}D_{N}(u,u)\geq(9/10)\cdot Tk\pi_{\star}$ (again, with high probability), we have that $\|M-(P+(k-1)\Pi)\|\leq\epsilon$.
8 Acknowledgements
All authors were partially supported by the Defence Science and Technology Laboratory.
[…]

### v2

@@ 5.1 An idealised setting
In Fig. 2 we present experimental results for a network topology corresponding to an undirected cycle of $6$ nodes. We can see that both transfer entropy and our methodology works well in recovering links in the network. In Fig. 2(d) we study the effect of a varying transmission length on network estimation. In particular, we record the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window. More precisely, we say that a link $\{i,j\}$ has been correctly identified by our methodology (resp. transfer entropy) if it appears in the top $m$ entries of the upper triangular part of the $\hat{L}_{\text{sym}}$ (resp. transfer entropy) matrix, where $m$ is the number of links in the network. [^1] We have adopted this simplistic inference procedure only to compare our estimator to transfer entropy. More complicated inference procedures, such as ones based on surrogate data approaches [25] or machine learning methods [6] might be needed in real world scenarios. We observe that a shorter transmission window makes estimation harder since it implies a larger number of nodes will be concurrently transmitting (i.e., a larger parameter $k$ in 1). We notice how our procedure slightly improves upon transfer entropy when the network is congested.
Further experimental results are reported in Figs. 3, 4, 5 and 6. We highlight that our methodology consistently outperforms transfer entropy whenever the network is congested. Transfer entropy and our estimate appear to work similarly well when $T\geq 15s$, with our estimator, however, giving more weight to pairs of nodes not connected by links: we believe this might be due to an underestimation of the value of $k$ defined in Section 3. A more refined approach for estimating $k$ might further improve our estimation procedure.
@@
Figure 2: Experimental results for a cyclic network topology of six nodes. Figure (a) displays the adjacency matrix of the network. Figure (b) the estimated matrix $\hat{L}_{\text{sym}}$ for a transmission window length $T=15s$. Figure (c) the transfer entropy matrix ($T=15s$). Figure (d) represents the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window. All results are averaged over $50$ independent simulations. A description of the experimental setup is provided at the beginning of Section 5.
@@
Figure 3: Experimental results for a network topology of eight nodes. See Fig. 2 for a description.
@@
Figure 4: Experimental results for a network topology where nodes are arranged in a three-by-three grid. See Fig. 2 for a description.
@@
Figure 5: Experimental results for a network topology where nodes are arranged in a four-by-two grid. See Fig. 2 for a description.
@@
Figure 6: Experimental results for a network topology of twenty nodes arranged in a path. See Fig. 2 for a description.
5.2 A more realistic setting
We now consider a more realistic scenario in which we do not have perfect information about the time series matrix TS. For simplicity, we will assume to know the number of nodes in the network, but not their position. To obtain information about packet transmission, we assume we are able to deploy a set of $s$ RF sensors in the environment. When a packet is transmitted from a node $u$, RF sensors positioned close to $u$ will be able to detect the corresponding power burst. We use the Log Distance path loss model in our ns-3 simulations: this implies that, the farther away from $u$ a sensor $\sigma$ is, the more attenuated the received signal by the sensor.
We again divide the time into small intervals $\{1,\dots,T^{\prime}\}$ (we assume a power burst is detected by at least a sensor at each interval, and discard intervals in which no power burst is detected); we can then construct a matrix $PD\in\mathbb{R}^{m\times T^{\prime}}$, where $PD(\sigma,t)$ represents the power detected at time $t$ by sensor $\sigma$. Suppose that, for a pair of time intervals $t,t^{\prime}$, the two vectors $PD(\cdot,t)$ and $PD(\cdot,t^{\prime})$ are somewhat similar: this indicates that the same node might be responsible for the detected power burst. This observation suggests a natural way to discern if the same node is transmitting during different time intervals: partition the vectors $\{PD(\cdot,t)\}_{t=1}^{T^{\prime}}$ using a clustering technique such as $k$-means. Time intervals assigned to the same cluster corresponds to power bursts we associate to the same node. In this way it is possible to construct a matrix which, hopefully, approximates the time series matrix TS (since we do not assume any information about the identity of the nodes, such as their position, this is only possible up to a relabelling of the nodes).
We perform experiments using the same ns-3 experimental setup discussed at beginning of the section, but instead of fixing the network topology we place $n$ transmitters and $s$ RF sensors at random in a rectangular box. For example, in Fig. 7 we have randomly placed $25$ transmitters and $100$ sensors in a square box of side $120m$. Only transmitters placed at most $50m$ apart are connected by a link in the network. This gives rise to a network topology with adjacency matrix displayed in Fig. 7(d). We simulate the network behaviour with ns-3 exchanging $500$ messages between randomly chosen pair of nodes. As explained above, we use $k$-means clustering (with $k=n$) to group together power bursts measured by different sensors in the same interval. Since we do not have any information about the identity of the transmitters (recall, we do not assume any information about the location of the sensors or transmitters), to evaluate how well measured power bursts are clustered together, we label each cluster computed by $k$-means according to the transmitter that corresponds to the largest number of bursts in the cluster. We display in Fig. 7(b) the percentage of power bursts detected that are associated to the correct transmitter for different lengths of the transmission window. The results are averaged over $20$ random realisations in which we fix the same network topology but randomly vary the position of the sensors. As expected, a larger transmission window results in a higher accuracy thanks to a lower number of concurrent transmissions.
We then compute our estimator $\hat{L}_{\text{sym}}$ together with the transfer entropy matrix $M_{\text{TE}}$. Examples of these two matrices for a particular random realisation are shown in Fig. 7 (e) and (f). Notice how these two matrices appear quite similar. In Fig. 7(c) we compare the performances of our method with transfer entropy. Again, we measured the number of times the top $m$ entries in the two matrices correspond to links in the network, where $m$ is the number of links in the network. Both methodologies appear to perform reasonably well for $T\geq 10$, with our estimation procedure slightly but consistently improving upon transfer entropy.
We repeat the same experiments with a bigger network: we randomly place $50$ transmitters and $150$ sensors in a rectangular box of $340m\times 200m$. We simulate the network with ns-3 exchanging now $2000$ messages between randomly chosen pairs. We perform $5$ random realisations, randomly placing both transmitters and sensors. Experimental results are shown in Fig. 8. We can observe a degradation of the performances, both in the proportion of identified transmitters (see Fig. 8(b)) and in the proportion of links of the network that corresponds to the top $m$ entries in both transfer entropy and $\hat{L}_{\text{sym}}$ matrix (see Fig. 8(c)). The first observation suggests that, in large networks, a simple application of $k$-means might not be enough to identify the correct transmitter for most power bursts; future work might want to exploit potential information about the position of sensors or different source separation procedures such as using independent component analysis [6]. The decrease in performance in correctly identifying links is not unexpected: the number of possible links increases quadratically in the number of transmitters while the number of actual links increases, roughly, linearly. Nevertheless, to help us better understand the reasons behind this decrease in performance, let us study closely Fig. 8 (f), which compares the top entries of $\hat{L}_{\text{sym}}$ against the actual links in the network: first of all, notice how the main “backbone” of the network is essentially preserved. As suggested by our theoretical analysis, our method does not necessarily preserve individual links but rather structural properties of the network. Furthermore, false positive links typically connect nodes that are only two hops apart in the network. This is not surprising since if there exist links $\{u,v\}$ and $\{v,w\}$, the time series corresponding to $u$ and $w$ will be correlated. This suggests it might be possible to reduce this error by, for example, combining information about time series with different interval lengths. We leave this as future work.
Figure 7: Experimental results for a random network constructed by placing $25$ transmitters and $100$ RF sensors at random in a $120m\times 120m$ square box. Figure (a) displays the position of the transmitters (black circles) and of the sensors (red crosses) for a particular random realisation, together with the resulting network topology. Figure (b) shows the number of power bursts correctly associated to the corresponding transmitters for a varying transmission window length (we display the average number over 20 random realisations, together with a 95% confidence interval). Figure (c) represents the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window (averaged over 20 random realisations). Figure (d) displays the adjacency matrix of the network. Figure (e) the estimated matrix $\hat{L}_{\text{sym}}$ for a transmission window length $T=10s$. Figure (f) the transfer entropy matrix ($T=10s$).
Figure 8: Experimental results for a random network constructed by placing $50$ transmitters and $150$ RF sensors at random in a $340m\times 200m$ rectangular box. Figure (a) displays the adjacency matrix of the network for a random realisation. Figure (b) an estimated matrix $\hat{L}_{\text{sym}}$ for a transmission window length $T=60s$. Figure (c) the corresponding transfer entropy matrix ($T=60s$). Figure (d) shows the number of power bursts correctly associated to the corresponding transmitters for a varying transmission window length (we display the average number over 5 random realisations together with a 95% confidence interval). Figure (e) represents the proportion of correctly identified links by our methodology and transfer entropy, for a varying length of the transmission window (averaged over 5 random realisations). Figure (f) displays (for a specific random realisation with window length $T=60s$) the correctly identified links by $\hat{L}_{\text{sym}}$ in green, and false positives in red; their thickness is proportional to the corresponding entry of $\hat{L}_{\text{sym}}$. Dashed black lines correspond to links in the network that do not appear among the top entries of $\hat{L}_{\text{sym}}$.
@@ 6 CONCLUSION AND FUTURE WORK
In this work, we have introduced a novel technique for estimating the topological structure of wireless networks. We demonstrated that our approach provably recovers information about the network topology under a simplified Markovian model of network dynamics. Additionally, our experimental results showed that the proposed method is competitive with transfer entropy in realistic network simulations, consistently outperforming it under conditions of network congestion.
@@
Finally, measures such as transfer entropy and Granger causality are often used as features in more sophisticated estimation procedures. For instance, Testi and Giorgetti [6] utilise these measures to train a neural network model. It would be interesting to explore the integration of our techniques into machine learning frameworks for network inference.
@@ 7 PROOFS
In this section we provide a proof of our main theoretical result, 1. Informally, we will show that, for $T$ large enough, $M(u,v)\approx P(u,v)-(k-1)\pi(v)$ and $N(v)\approx T\cdot k\cdot\pi(v)$. Therefore, by computing $M$ and $N$, we are able to approximate $P$. We start by showing that $N(v)$ is approximately a multiple of the stationary distribution at $v$. For technical reason, we also need a bound on the following quantities ($u,v\in V$):
$\widetilde{N}(u,v)=\sum_{t=1}^{T}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\},$ (2)
$Q(u,v)=\sum_{t=1}^{T}\sum_{i,i^{\prime}}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(i^{\prime})}=v\}.$ (3)
We will use the following recent matrix Bernstein inequality for Markov chains [22].
@@
We use 2 to bound the probability $N$, $\tilde{N}$, and $Q$ deviate from their expectation.
@@
Let $\epsilon\in(0,1/2),\delta\in(0,1)$. Let $T>Ck^{2}(\pi^{\star}/\pi_{\star})\ln(kn/\delta)/(\epsilon^{2}\lambda(P)\pi_{\star})$ for some large enough universal constant $C>0$. Let $D_{N}$ be the $n\times n$ diagonal matrix such that $D_{N}(u,u)=N(u)$. Then, with probability at least $1-\delta$, for all $v\in V$, the following holds:
@@
$\left\|Q-Tk(k-1)\pi\pi^{T}-D_{N}\right\|\leq Tk^{2}\pi^{\star}.$
@@
$\mathop{\mathbb{E}}[N(v)]$ $=\sum_{i=1}^{k}\sum_{t=1}^{T}\mathop{\mathbb{P}}[X^{(j)}_{t}=v]=kT\pi(v),$
@@
$\mathop{\mathbb{E}}[\widetilde{N}(u,v)]$ $=\sum_{t=1}^{T}\sum_{i\neq j}\mathop{\mathbb{E}}[\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}]$
@@
Notice that $\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}]=\sum_{i\neq j}\sum_{t=1}^{T}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})$. Clearly, $\|g_{ij}((u,v))\|\leq 2$ for all $(u,v)\in V\times V$. Furthermore, simple calculations show that $\|\mathop{\mathbb{E}}_{#1}[g_{ij}((u,v))^{2}]\leq 2\pi^{\star}\|$. Therefore, with probability at least $\delta/k^{2}$, Theorem 2 implies that $\left\|\sum_{t=1}^{T}g_{ij}(X_{t}^{(i)},X_{t}^{(j)})\right\|\leq Tk\pi_{\star}$. A union bound over all $1\leq i\neq j\leq k$ yields the second statement.
Finally, the third inequality follows simply from the first two by noticing that $Q(u,v)=\widetilde{N}(u,v)+D_{N}(u,v)$. ∎
We now prove concentration for $M$. We apply similar techniques as in Kontorovich and Wolfer [13]. We will need the following concentration inequality for matrix martingales, which is due to Tropp [23].
@@
where with $\mathop{\mathbb{E}}_{#1}[Y_{k}]$ we denote the conditional expectation of $Y_{k}$ w.r.t. “the past”.
@@
$\mathop{\mathbb{E}}_{#1}[\mathbbm{1}\{X_{t-1}^{(i)}=u\}\mathbbm{1}\{X_{t}^{(j)}=v\}]$
$\quad=\mathbbm{1}\{X_{t-1}^{(i)}=u\}\sum_{w\in V}\mathbbm{1}\{X_{t-1}^{(j)}=w\}P(w,v).$
@@
Let $\widetilde{N}(u,v)=\sum_{t=1}^{T}\sum_{i\neq j}\mathbbm{1}\{X_{t}^{(i)}=u,X_{t}^{(j)}=v\}$. Then, for any $u,v\in V$,
@@
$\|\widetilde{N}P-k(k-1)T\pi\pi^{T}\|$ $<\|\mathop{\mathbb{E}}[\widetilde{N}]P-k(k-1)T\pi\pi^{T}\|$
$\quad+\|(\widetilde{N}-\mathop{\mathbb{E}}[\widetilde{N}])P\|$
@@
$\leq\max_{v\in V}\left\{\left|\sum_{i\colon X^{(i)}_{t}=v}(1-P(X^{(i)}_{t-1},v))\right|\right.$
$\quad+k\cdot\sum_{i=1}^{k}\left.\left|\sum_{j\colon X^{(j)}_{t}=v}(1-P(X^{(j)}_{t-1},v))\right|\right\}$
@@
$\sum_{v\in V}\left|\left(\mathbbm{1}\{X_{t}^{(i)}=v\}-P(u,v)\right)\right|$
$\quad=1-P(u,X_{t}^{(i)})+\sum_{w\neq X_{t}^{(i)}}P(u,w)$
$\quad=2(1-P(u,X_{t}^{(i)})).$
@@
$\leq\max_{u\in V}\left\{2\sum_{i=1}^{k}(1-P(u,X_{t}^{(i)}))\right.$
$\quad+\left.\sum_{i=1}^{k}\left|\sum_{j\neq i}2(1-P(X^{(j)}_{t-1},v))\right|\right\}$
@@
$\|Y_{t}\|\leq\sqrt{\|Y_{t}\|_{1}\cdot\|Y_{t}\|_{\infty}}\leq\sqrt{2}k^{2}.$
@@
$\qquad\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i^{\prime})}=v\}-P(w,v)\right)$
@@
$\qquad\cdot\left(\mathbbm{1}\{X_{j}^{(i)}=v\}-P(w,v)\right)\Bigr{]}$
@@
By Lemma 3, with high probability, we have that $N(u)\leq 2kT\pi(u)$ for any $u\in V$. Therefore,
@@
We have proved that $\|N-(D_{N}+\widetilde{N})P\|=\|D_{N}(M-P-\widetilde{M}P))\|\leq(1/\sqrt{2})\cdot Tk\pi_{\star}\epsilon$ with high probability. By (5), we have that $\|D_{N}(M-P-(k-1)\Pi)\|\leq(5/6)\cdot Tk\pi_{\star}\epsilon$ with high probability. Since $\min_{u}D_{N}(u,u)\geq(9/10)\cdot Tk\pi_{\star}$ (again, with high probability), we have that $\|M-(P+(k-1)\Pi)\|\leq\epsilon$. This ends the proof of 1.